### v1

# Classification of foliations via disintegration maps

Abstract.
In this paper, we present a novel approach for analyzing the relationship between the supports of conditional measures and their geometric arrangement in Wasserstein space via the disintegration map. Our method establishes criteria to determine when such conditional measures arise from a metric measure foliation. Additionally, we provide a concrete example demonstrating how this framework can be applied to study deformations (or perturbations) of disintegration-induced foliations.
1. Introduction
The disintegration of measures offers a powerful framework for analyzing local and global probabilistic structures by decomposing a measure into conditional probabilities. Roughly speaking, given $X,Y$ measurable spaces, $\mu$ a measure on $X$, and a measurable map $\pi:X\to Y$, a disintegration allows us to break $\mu$ into a family of conditional measures $\{\mu_{y}\}_{y\in Y}$ on $X$, depending measurably on $y$, such that $\mu$ can be reconstructed by integrating $\mu_{y}$ against $\pi_{*}\mu$, the push-forward of $\mu$ by $\pi$. This decomposition is crucial in several areas, such as Probability Theory, Ergodic Theory, and Geometric Measure Theory, where it provides a rigorous way to handle conditional distributions.
Disintegration theorems typically require topological assumptions to ensure existence and uniqueness of such decompositions, making them a powerful tool to define measures on fibered spaces or under projections, for instance. Although significant progress has been made in establishing the existence of conditional measures under several different hypothesis, most of the existing approaches neglect the geometric aspects of the underlying space.
In a recent paper [PR25], the second and third authors used the intrinsic structures of probability spaces to tackle the disintegration of measures from a perspective which takes into accounting statistical and geometrical properties of dynamical systems. In particular, it is introduced the notion of disintegration maps, which are essentially functions that assign to each point $y\in Y$ its corresponding conditional measure $\mu_{y}$, providing a structured way to analyze a disintegration. We emphasize that these objects can trace geodesics in the space of probability measures, which correspond to the solutions of a time-evolving transport problem, and exhibit certain rigidity phenomena: if one of the endpoints measure in the geodesic path is absolutely continuous, then so are all. Therefore, it becomes evident that certain geometric properties of the base space control the regularity of the conditional measures.
In this paper, we further explore disintegration maps by defining a notion of derivative for these objects. Our notion of derivative resembles a metric derivative, comparing in a appropriated way the distance between conditional measures and the distance between the supports of such measures. We remark that, in order to define this notion of derivative we do not require any differentiable structure on the base space. An interesting case arises in the context of Riemannian submersions, for example, since the distance between fibers equals the distance between base points. In this context, if each conditional measure has full support on its corresponding fiber, we show that the distance among conditional measures and the distance between base points are the same, and this isometry is captured by the derivative. We introduce a novel dense energy functional, similarly to what is done in geometric analysis. This notion is suited to quantify the relationship between these distances more accurately. We show that the minimization of this functional detect and characterize highly ordered geometric structures in the base space. More specifically, we show that the minimum of this energy functional corresponds to a metric measure foliation. That is, a foliation of the base space whose leaves are parallel to each other and where the distance between conditional measures on these leaves equals the distance between the leaves themselves. Moreover, this functional exhibits sensitivity to small deformations of the foliation, capturing fine geometric structures. We provide an explicit example in which we assess the sensitivity of the energy functional by inducing deformations in a metric measure foliation.
The main contributions of this paper are presented in Section 4, where we introduce a novel notion of derivative for disintegration maps (Definition 4.1). We then construct an energy functional designed to quantify the relationship between the Wasserstein distance of conditional measures and the distance between their supports (Definition 4.2). Our main result is:
Theorem 1.1.
The minimum of $\mathcal{E}(f)$, the energy of the disintegration map, is equals to one, if and only if, the supports of conditional measures form a metric measure foliation.
The precise statement shall be given in Theorem A.
The paper is organized as follows. Section 2 outlines the theoretical framework and necessary definitions. Section 3 reviews essential background on disintegration maps. Finally, Section 4 presents our core results, including a detailed example exploring the sensitivity of the energy functional to deformations of the leaves that support the conditional measures.
2. Theoretical framework
Let $X$, $Y$ be locally compact complete separable metric spaces. Consider $\pi:X\to Y$ a Borel map, and $\mu\in\mathcal{M}_{+}(X)$, where $\mathcal{M}_{+}(X)$ is the set of all finite positive Radon measures on $X$. Define $\nu:=\pi_{*}\mu$ in $\mathcal{M}_{+}(Y)$. Then, by [PR25, Theorem A] there exist measures $\mu_{y}\in\mathcal{M}_{+}(X)$, such that,
  1. (1)
    $y\mapsto\mu_{y}$ is a Borel map for $\nu$-almost every $y\in Y$;
  2. (2)
    $\mu_{y}$ is a probability on $X$ for $\nu$-almost every $y\in Y$;
  3. (3)
    $\mu_{y}$ is concentrated on $\pi^{-1}(y)$ for $\nu$-almost every $y\in Y$;
  4. (4)
    $\mu(A)=\int_{Y}\mu_{y}(A)~d\nu(y)$ for every Borel subset $A$ of $X$.
The family $\{\mu_{y}\}_{y\in Y}$ is called the disintegration of $\mu$ with respect to $\nu$.
in order to examine this family of probability measures, we endow $\mathscr{P}(X)$, the space of probability measures on $X$, with a metric structure. The framework of Wasserstein spaces provides a natural geometric setting for such investigations. The Wasserstein space of order $p$, for $p\in[1,\infty)$, is the set
$\mathscr{P}_{p}(X):=\Big{\{}\mu\in\mathscr{P}(X):\int d(x,\tilde{x})^{p}\mu(dx)<+\infty\Big{\}}$
$\tilde{x}\in X$ arbitrary, endowed with the Wasserstein distance
$W_{p}(\mu,\nu):=\Big{(}\inf\limits_{\gamma\in\Pi(\mu,\nu)}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})\Big{)}^{\frac{1}{p}}.$
for $\mu$ and $\nu$ probability measures on $X$, where $\Pi(\mu,\nu)$ is the set of measures $\gamma\in\mathscr{P}(X\times X)$ with marginals $\mu$ and $\nu$.
Given a disintegration of $\mu$ with respect to $\nu$, we call disintegration map a function assigns to each $y\in Y$ a conditional measure $\mu_{y}\in\mathscr{P}_{p}(X)$.
Definition 2.1.
[PR25, Definition 4.1] Let $X$, $Y$ be locally compact complete separable metric spaces. Consider $\pi:X\to Y$ a Borel map, $\mu\in\mathcal{M}_{+}(X)$, and $\nu:=\pi_{*}\mu$. Given $\{\mu_{y}\}_{y\in Y}$ a disintegration of $\mu$ with respect to $\nu$, we define the disintegration map by
$f:D\subset Y$ $\to(\mathscr{P}_{p}(X),W_{p})$
$y$ $\mapsto\mu_{y},$
such that $\mu(A)=\int_{Y}f(y)(A)~d\nu(y)$, for every Borel subset $A$ of $X$. In order to make clear which measures are associated with the disintegration map, we say that “$f$ is a disintegration map of $\mu$ with respect to $\nu$”.
To give an example, consider $X=[0,1]\times[0,1]$ the unit square, $Y=[0,1]$, $\mu=\text{Leb}^{2}$ the two-dimensional Lebesgue measure on $X$, $\pi(x,y)=x$ the projection onto the first coordinate, and $\nu=\text{Leb}^{1}$ the one-dimensional Lebesgue measure on $Y$. The disintegration $\{\mu_{x}\}_{x\in[0,1]}$ of $\mu$ with respect to $\nu$ is given by:
$\mu_{x}=\text{Leb}^{1}\text{ on }\{x\}\times[0,1]\quad\text{(vertical line segment at }x\text{)}$
Explicitly, for any measurable $A\subseteq X$:
$\mu(A)=\int_{[0,1]}\mu_{x}(A)\,d\nu(x)=\int_{0}^{1}\text{Length}\big{(}A\cap(\{x\}\times[0,1])\big{)}\,dx$
where $\text{Length}(A\cap(\{x\}\times[0,1])):=\text{Leb}^{1}(y\in[0,1]~:~(x,y)\in A)$. In fact, note that the map $x\mapsto\mu_{x}(A)=\text{Length}(A\cap\{x\}\times[0,1])$ is measurable, for any $B\subseteq[0,1]$,
$\mu(A\cap\pi^{-1}(B))=\int_{B}\text{Length}\big{(}A\cap\{x\}\times[0,1]\big{)}\,dx$
and $\mu_{x}$ is supported on $\pi^{-1}(x)=\{x\}\times[0,1]$, satisfying the disintegration conditions.
Figure 1. Disintegration of $\text{Leb}^{2}$ with respect to $\text{Leb}^{1}$: the red line $\{x\}\times[0,1]$ carries $\mu_{x}=\text{Leb}^{1}$
In this case, the disintegration map is given by
$f:[0,1]$ $\to(\mathscr{P}_{p}([0,1]\times[0,1]),W_{p})$
$x$ $\mapsto\mu_{x},$
where $\mu_{x}$ is $\text{Leb}^{1}$ on $\{x\}\times[0,1]$. This is a basic example that fits into the broader framework of metric foliations, which is the primary focus of our study.
Let $(X,d)$ be a metric space, and $\mathcal{F}$ a partition of $X$ into closed subsets. We call $\mathcal{F}$ a foliation of $X$ and the elements of this partition are called leaves. If for every $F,F^{\prime}\in\mathcal{F}$ and every $x\in F$, $d(F,F^{\prime})=d(x,F^{\prime})$, where $d(F,F^{\prime})=\inf\{d(x,x^{\prime}):x\in F,x^{\prime}\in F^{\prime}\}$ and $d(x,F^{\prime})=d(\{x\},F^{\prime})$, then $\mathcal{F}$ is called a metric foliation. In case that each leaf is bounded, we say that $\mathcal{F}$ is bounded. Given a metric foliation $\mathcal{F}$ of $X$, define the equivalence relation:
$x\sim x^{\prime}~\iff~\exists~F\in\mathcal{F}~\text{such that}~x,x^{\prime}\in F.$ (1)
Consider the $X^{*}:=X/\sim$ the set of equivalence classes under (1) and the projection $\pi:X\to X^{*}$ onto $X^{*}$. We call $X^{*}$ the quotient space and $\pi$ the quotient map. Define a distance function $d^{*}$ on $X^{*}$ as
$d^{*}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))$ (2)
for $y,y^{\prime}\in X^{*}$. Note that $\pi$ is a submetry: $\pi(B(x,r))=B(\pi(x),r)$, where $B(x,r)$ is a ball centred at $x$ with radius $r$. In fact,
$B(\pi(x),r)=$ $~\{y\in X^{*}:d^{*}(y,\pi(x))<r\}$
$=$ $~\{y\in X^{*}:d(\pi^{-1}(y),\pi^{-1}(\pi(x)))<r\}$
$=$ $~\pi(B(x,r)).$
Therefore, $\pi$ is 1-Lipschitz. In this notation, we define a metric measure foliation.
Definition 2.2.
Let $X$ be locally compact complete separable metric space, and $\mathcal{F}$ be a metric foliation of $X$. Consider $\mu\in\mathscr{M}_{+}(X)$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$. We call $\mathcal{F}$ a $p$-metric measure foliation, $p\in[1,\infty)$, if $\pi_{*}\mu$ is locally finite Borel measure on $X^{*}$, and
$W_{p}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})$ (3)
for any $y,y^{\prime}\in Y$, where $\{\mu_{y}\}_{y\in Y}$ is a disintegration of $\mu$ with respect to $\pi_{*}\mu$.
Remark 2.3.
The general definition for a metric measure foliation requires $p=2$ and (3) only almost everywhere (see [GKMS18]). However, our definition is more restrictive, requiring (3) everywhere.
Note that, in this case, the disintegration map is an isometry. This is the most important characteristic of metric measure foliation for us. Although the definition holds for $p\in[1,\infty)$, our case of interest is $p=2$, since $\mathscr{P}_{2}(X)$ has a rich geometric structure [GRS25]. Moreover, if $\mathcal{F}$ is a 2-metric measure foliation, then
$W_{q}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))$
for any $q\in[1,\infty)$ and any $y,y^{\prime}\in X^{*}$, where $\{\mu_{y}\}_{y\in Y}$ is the disintegration of $\mu$ with respect to $\pi_{*}\mu$ [Kaz22, Proposition 3.11]. Henceforth, the term “metric measure foliation” will refer to what we have defined as a “2-metric measure foliation”.
A very important example of metric measure foliation is related to the action of isometry group [GR23]. Let $(X,d,\mu)$ be a metric measure space and $G$ a compact topological group. Let
$G\times X\ni(g,x)\mapsto gx\in X$
be an isometric action of $G$ on $X$. Suppose this action is metric measure isomorphic. That is, for every $g\in G$ the map $X\ni x\mapsto gx\in X$ is an isometry preserving the measure $\mu$. Consider $[x]$ the $G$-orbit of a point $x\in X$ and the quotient space $X/G$ endowed with the distance
$d_{X/G}([x],[x^{\prime}])=\inf_{g,g^{\prime}\in G}d(gx,g^{\prime}x^{\prime}).$
Consider $\pi:X\to X/G$ the projection map, that is, $\pi$ is given by $x\mapsto[x]$. The family $\mathcal{F}:=\{\pi^{-1}(y):y\in X/G\}$ is a metric measure foliation on $X$.
Other interesting examples arise from Riemannian submersions of weighted Riemannian manifolds [GKMS18], for instance.
3. Preliminaries on disintegration maps
In this section, we discuss key properties of disintegration maps, as introduced in Definition 2.1 and developed in [PR25]. To start our discussion, suppose $X=Y=\mathbb{R}$, and let $\gamma$ be a probability measure supported on the graph of a measurable function $\pi:X\to Y$, i.e., $\gamma=(\mathrm{id},\pi)_{*}\mu$, where $\mu$ its a probability measure on $X$ and $(\mathrm{id},\pi)_{*}\mu$ is the pushforward of $\mu$ under the map $x\mapsto(x,\pi(x))$. Then, the disintegration of $\gamma$ with respect to $\mu$ is given by the family of Dirac measures:
$\gamma_{x}=\delta_{\pi(x)}.$
In fact, for any measurable $A\subseteq X\times Y$, we have
$\gamma(A)=\mu(\{x\in X:(x,\pi(x))\in A\}).$
By definition of the disintegration
$\gamma(A)=\int_{X}\gamma_{x}(A_{x})\,d\mu(x)=\int_{X}\delta_{\pi(x)}(A_{x})\,d\mu(x).$
where $A_{x}=\{y\in Y:(x,y)\in A\}$. Then,
$\delta_{\pi(x)}(A_{x})=\mathbbm{1}_{A}(x,\pi(x)),$
and thus:
$\gamma(A)=\int_{X}\mathbbm{1}_{A}(x,\pi(x))\,d\mu(x),$
which matches the definition of $\gamma$. This shows that when $\gamma$ is supported on the graph of a function $\pi$, its disintegration is given by $\gamma_{x}=\delta_{\pi(x)}$. This is a fundamental example in Optimal Transport and conditional probability, where deterministic couplings lead to Dirac disintegrations. Observe that when $\pi=\mathrm{id}$, we have a natural isometric immersion of $X$ into $\mathscr{P}_{p}(X)$, namely the disintegration map $x\mapsto\delta_{x}$.
In general, we aim to investigate under what conditions geometric information about $X$ can be derived via disintegration — specifically, through the properties of disintegration maps. A key starting point for this analysis is understanding the continuity of these maps. When referring to continuity of the disintegration map, we mean continuity with respect to weak convergence on $\mathscr{P}_{p}(X)$. In [PR25] some results are presented regarding the continuity of disintegration maps, namely:
  1. (1)
    [PR25, Propostion 5.2] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$. If $\pi:X\to Y$ is a Borel map such that $\nu:=\pi_{*}\mu$ is a Borel measure, then, the disintegration map $f:Y\to(\mathscr{P}_{p}(X),W_{p})$ of $\mu$ with respect to $\nu$ is nearly continuous, for $p\in[1,\infty)$. A map $f$ on an metric space $(Y,\nu)$ is called nearly continuous if, for each $\varepsilon>0$ there exist $\mathcal{K}\subset Y$ closed with $\nu(Y\backslash\mathcal{K})<\varepsilon$ such that $f$ restricted to $\mathcal{K}$ is continuous.
  2. (2)
    [PR25, Propostion 5.4] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$. If the Borel map $\pi:X\to Y$ is bijective and $\pi^{-1}$ continuous, then the disintegration map $f$ of $\mu$ with respect to $\pi_{*}\mu$ is continuous.
  3. (3)
    [PR25, Propostion 5.8] Let $X$ be locally compact complete separable metric space. Considering a metric foliation $\mathcal{F}$ of $X$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$, if there exists a metric measure foliation of $X$, then the disintegration map of $\mu$ with respect to $\nu:=\pi_{*}\mu$ is continuous.
The continuity of disintegration maps enables us to relate paths in the chosen indexing metric space to corresponding paths of measures in Wasserstein space. For instance, in the case of the disintegration map $x\mapsto\delta_{x}$, there is a direct correspondence between a path in $X$ and the induced path of Dirac measures. In the context of a metric measure foliation, this relationship becomes more nuanced. Here, transversal paths (those crossing the leaves) must satisfy an additional condition concerning the distance between the supports of the conditional measures along the path — a consequence of the parallelism of the leaves.
In the next section we present an approach that allows us to study the relation between the relative position of the elements of the family of supports and the relative position of the conditional measures in the Wasserstein space, via a disintegration map.
4. On foliations and disintegration maps
Motivated by the concept of the metric derivative and the interplay between the relative positions of supports and conditional measures, we introduce a notion of derivative for disintegration maps to quantify how these measures evolve along paths.
Definition 4.1.
Let $X,Y$ be locally compact complete separable metric spaces. Consider a measure $\mu\in\mathcal{M}_{+}(X)$, $\pi:X\to Y$ a Borel map, $\nu=\pi_{*}\mu$, and $f$ the disintegration map of $\mu$ with respect to $\nu$. We define a notion of derivative of $f$ by
$|\nabla f(y)|_{p}:=\lim\limits_{\varepsilon\to 0}\sup_{y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)}\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))}$
for $p\in[1,\infty)$, where $\mu_{y^{\prime}}=f(y^{\prime})$ and $\mu_{y^{\prime\prime}}=f(y^{\prime\prime})$ are conditional measures and $B_{\varepsilon}(y)$ is a ball centered in $y\in Y$ with radius $\varepsilon$.
Let $d_{Y}(y^{\prime},y^{\prime\prime})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$ denote the distance between the fibers over $y^{\prime}$ and $y^{\prime\prime}$. In the case of a metric measure foliation, the $p$-Wasserstein distance between conditional measures satisfies $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d_{Y}(y^{\prime},y^{\prime\prime})$, which implies $|\nabla f(y)|_{p}=1$. This derivative concept thus serves as a tool to characterize when $\{\pi^{-1}(y)\}$ forms a metric measure foliation. To further analyze the connection between the distances of conditional measure supports and their Wasserstein distance, paralleling techniques in geometric analysis, we introduce a dense energy functional designed to quantify this relationship.
Definition 4.2.
Let $X,Y$ be locally compact complete separable metric spaces. Consider a measure $\mu\in\mathcal{M}_{+}(X)$, a Borel map $\pi:X\to Y$, and the disintegration map $f$ of $\mu$ with respect to $\nu=\pi_{*}\mu$. The $p$-energy of $f$ is given by
$\mathcal{E}_{p}(f):=\|\nabla f\|_{\infty,p}=\sup_{y\in Y}|\nabla f(y)|_{p}.$
Note that $|\nabla f|_{p}\geq 1$ everywhere. In fact, since $\mu_{y^{\prime}}$ and $\mu_{y^{\prime\prime}}$ are supported on $\pi^{-1}(y^{\prime})$ and $\pi^{-1}(y^{\prime\prime})$, respectively, we have
$W_{p}^{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}}):=$ $\inf\limits_{\gamma\in\Pi(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})$
$\geq$ $~d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))^{p}.$
Moreover, note that $\min_{f}\mathcal{E}_{p}(f)=1$ if, and only if, $\|\nabla f\|_{\infty,p}=1$, which in turn hold true if, and only if $|\nabla f|_{p}=1$. In this case we have an isometry $W_{p}=d_{Y}$. Under some further hypotheses, we can relate the functional $\mathcal{E}_{p}(f)$ to the metric measure foliation case.
Theorem A.
Let $(X,d)$ be a locally compact complete separable geodesic space, $Y$ a compact complete separable metric space, and $\pi:X\to Y$ a Borel map. Given $\mu\in\mathcal{M}_{+}(X)$, let $f_{\mu,\pi}$ be a disintegration of $\mu$ with respect to $\nu:=\pi_{*}\mu$. Suppose that the conditional measures $\mu_{y}$ have full support, i.e., $\operatorname{supp}(\mu_{y})=\pi^{-1}(y)$. Then, $\mathcal{E}_{p}(f_{\mu,\pi})=1$ if, and only if, $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$. Therefore, the energy-minimizing disintegration maps correspond to a metric measure foliation.
Proof.
[…]
$u(y^{\prime},y^{\prime\prime}):=\begin{cases}\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{d_{Y}(y^{\prime},y^{\prime\prime})}&~\text{,if}~~y^{\prime}\neq y^{\prime\prime}\\ 1&~\text{,if}~~y^{\prime}=y^{\prime\prime}\end{cases}$
Note that $u\geq 1$. By the hypothesis,
$\lim\limits_{\varepsilon\to 0}\sup\limits_{y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)}u(y^{\prime},y^{\prime\prime})=u(y,y)$
for every $y\in Y$. Then, $u$ is continuous at $(y,y)$ for every $y\in Y$. Moreover, $u$ is continuous everywhere else. In fact, since $\min_{f_{\mu,\pi}}\mathcal{E}_{p}(f_{\mu,\pi})=1$, we have $|\nabla f(y)|p=1$ for every $y\in Y$, and the disintegration map $f$ is continuous. Thus, it implies the continuity of $u$ everywhere else.
Since $Y$ compact, $Y\times Y$ is compact and by compactness, we have uniform continuity, that is, for all $\varepsilon>0$ there exists $\delta>0$ such that,
$d_{Y\times Y}((y,y^{\prime}),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$
implies $|u(y,y^{\prime})-u(\tilde{y},\tilde{y}^{\prime})|\leq\varepsilon$, for every ($y~)(y,y^{\prime}$),($\tilde$$\tilde{y}^{\prime}$), where
[…]
Note that $|u(y,y)-u(\tilde{y},\tilde{y}^{\prime})|=u(\tilde{y},\tilde{y}^{\prime})-1$. Then, setting $y=\tilde{y}$, and if $d_{Y}((\tilde{y},\tilde{y}^{\prime}),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$, we get $d_{Y\times Y}((y,y),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$, Thus,
$|u(y,y)-u(\tilde{y},\tilde{y}^{\prime})|=u(\tilde{y},\tilde{y}^{\prime})-1\leq\varepsilon,$
[…]
$\sum_{k=0}^{n-1}d(x_{k},x_{k+1})=d(x,x^{\prime}).$
Then,
[…]
$\leq$ $\sum_{k=0}^{n-1}d(\pi^{-1}(\pi(x_{k})),\pi^{-1}(\pi(x_{k+1})))$
$\leq$ $(1+\varepsilon)\sum_{k=0}^{n-1}d(\pi^{-1}(\pi(x_{k})),\pi^{-1}(\pi(x_{k+1})))$
$\leq$ $(1+\varepsilon)\sum_{k=0}^{n-1}d(x_{k},x_{k+1})$
$=$ $(1+\varepsilon)d(x,x^{\prime})$
[…]
Let $\gamma$ be the optimal transport plan from $\mu_{y}$ to $\mu_{y^{\prime}}$, and consider, for all $(x,x^{\prime})\in\operatorname{supp}(\gamma)$, $\varphi(x)=d(x,\pi^{-1}(y^{\prime}))>d(x,x^{\prime})$. Then,
$\int\varphi(x)d\gamma=\int\varphi(x)d\mu_{y}=\int d(x,\pi^{-1}(y^{\prime}))d\mu_{y}>\int d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))d\mu_{y}$
that is a contradiction, since
$W_{1}(\mu_{\tilde{y}},\mu_{\tilde{y}^{\prime}}):=\inf\limits_{\gamma\in\Pi(\mu_{\tilde{y}},\mu_{\tilde{y}^{\prime}})}\int d(x_{1},x_{2})~d\gamma(x_{1},x_{2}).$
An analogous argument applies to the case $p\in(1,\infty)$.
Remark 4.3.
The conclusion of Theorem A does not hold if full support is not assumed. In fact, consider $X=[-1,1]\times\mathbb{R}$, $Y=[0,1]$, $\pi:X\to Y$ such that $\pi^{-1}(y)$ is a ellipse given by $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$, with major radius $y$. Consider a measure $\mu\in\mathcal{M}_{+}(X)$ such that the disintegration of $\mu$ with respect to $\nu$ is $\mu_{y}=\delta_{(0,\frac{y}{\lambda})}$. In this case,
$d_{Y}(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\left|\frac{y}{\lambda}-\frac{y^{\prime}}{\lambda}\right|=W_{p}(\mu_{y},\mu_{y^{\prime}})$
for every $y,y^{\prime}\in Y$. Therefore, we have isometry, but $\{\pi^{-1}(y)\}$ does not define a metric foliation. A graphical representation of the case $\lambda=1.5.$ is provided in Figure 2.
Figure 2. Illustration for the case in which the measures $\mu_{y}$ do not have full support: there is isometry between $W_{p}$ and $d_{Y}$, but $\{\pi^{-1}(y)\}$ (the ellipses) does not define a metric foliation.
Note that in the previous example, if $\lambda=1$, that is, the preimages are circles in $X$, then $\{\pi^{-1}(y)\}$ forms a metric foliation. In the context of Theorem A, we aim to demonstrate that the $p$-energy exhibits sensitivity to small perturbations of the foliation. To analyze this, we first evaluate the $1$-energy $\mathcal{E}_{1}$ for a metric measure foliation on $\mathbbm{R}^{2}$ with concentric circular leaves. We then generalize this analysis to cases where the leaves are slightly deformed into ellipses.
Example 4.4.
Consider $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}\leq R^{2}\}$, $R>0$, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed, and $Y=(0,R]$. Consider a foliation of $X$ into circles, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+x_{2}^{2}}=y$. To simplify the analysis, consider this setup in polar coordinates: consider $x_{1}=r\cos(\theta)$, $x_{2}=r\sin(\theta)$. Using Cartesian coordinates, an infinitesimal area element can be calculated as
$dA=dx_{1}~dx_{2}.$
To convert it to polar coordinates, the Jacobian determinant $J$ of the coordinate conversion formula has to be considered. In this case:
$J(r,\theta)=\begin{vmatrix}\frac{\partial x_{1}}{\partial r}&\frac{\partial x_{1}}{\partial\theta}\\ \frac{\partial x_{2}}{\partial r}&\frac{\partial x_{2}}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos(\theta)&-r\sin(\theta)\\ \sin(\theta)&r\cos(\theta)\end{vmatrix}=r.$
Then,
$dA=r~dr~d\theta.$
Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
$d\mu=\frac{dA}{A}=\frac{r~dr~d\theta}{\pi R^{2}}.$
Define $\nu:=\pi_{*}\mu$. Note that, give a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between circles of radii $a$ and $b$, normalized by $A$:
$\nu((a,b))=\frac{\pi b^{2}-\pi a^{2}}{\pi R^{2}}=\frac{b^{2}-a^{2}}{R^{2}}.$
In this way,
$d\nu(r)=\frac{2r}{R^{2}}~dr.$
Let $\{\mu_{y}\}_{y\in Y}$ be the disintegration of $\mu$ with respect to $\nu$, such that each $\mu_{y}$ is a uniformly distributed probability measure on $F_{y}$. Due to the symmetry of the circle, the arc length is proportional to $\theta$, and
$d\mu_{y}(r,\theta)=\frac{1}{2\pi}~\delta_{y}~d\theta.$
Note that
$\int_{Y}\mu_{y}~\nu=\int_{0}^{R}\int_{0}^{2\pi}\Big{(}\frac{1}{2\pi}\delta_{y}d\theta\Big{)}\frac{2r}{R^{2}}~dr=\frac{1}{\pi R^{2}}\iint r~dr~d\theta$
so we recover the measure $\mu$ by integrating the family $\mu_{y}$ with respect to $\nu$. Moreover, $d_{Y}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=|y-y^{\prime}|$. For simplicity, consider $W_{1}$. One way to write the 1-Wasserstein distance is
$W_{1}(\mu_{y},\mu_{y^{\prime}})=\inf_{T}\int d(x,T(x))~d\mu_{y}$
and since we have measures uniformly distributed and symmetry in the leaves, the function $T$ that minimizes this problem is such that, for every $x\in F_{y}$, $T(x)$ is a point in $F_{y^{\prime}}$ such that $d(x,T(x))=|y-y^{\prime}|$, that is, the transport occurs radially, orthogonally to the leaves. Therefore,
$W_{1}(\mu_{y},\mu_{y^{\prime}})=\int_{F_{y}}|y-y^{\prime}|~d\mu_{y}=|y-y^{\prime}|.$
Then, the disintegration map $\tilde{f}$ of $\mu$ with respect to $\nu$ is a isometry, and $\mathcal{E}_{1}(\tilde{f})=1$.
Similarly, consider the case $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}\leq R^{2}\}$, with $R>0$ and $\lambda=1+\varepsilon$, $\varepsilon$ positive and sufficiently small, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed, $Y=(0,R]$, and the foliation of $X$ is given into ellipses, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+\lambda^{2}x_{2}^{2}}=y$. Converting into polar coordinates: consider $x_{1}=r\cos(\theta)$, $x_{2}=\frac{r}{\lambda}\sin(\theta)$, and
$J(r,\theta)=\begin{vmatrix}\frac{\partial x_{1}}{\partial r}&\frac{\partial x_{1}}{\partial\theta}\\ \frac{\partial x_{2}}{\partial r}&\frac{\partial x_{2}}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos(\theta)&-r\sin(\theta)\\ \frac{1}{\lambda}\sin(\theta)&\frac{r}{\lambda}\cos(\theta)\end{vmatrix}=\frac{r}{\lambda}.$
Then,
$dA=\frac{r}{\lambda}~dr~d\theta.$
Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
$d\mu=\frac{dA}{A}=\frac{\frac{r}{\lambda}~dr~d\theta}{\frac{\pi R^{2}}{\lambda}}=\frac{r~dr~d\theta}{\pi R^{2}}.$
Define $\nu:=\pi_{*}\mu$. Note that, give a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between ellipses of major axis $a$ and $b$, normalized by $A$:
$\nu((a,b))=\frac{\frac{\pi b^{2}}{\lambda}-\frac{\pi a^{2}}{\lambda}}{\frac{\pi R^{2}}{\lambda}}=\frac{b^{2}-a^{2}}{R^{2}}.$
In this way,
$d\nu(r)=\frac{2r}{R^{2}}~dr.$
Let $\{\mu_{y}\}_{y\in Y}$ be the disintegration of $\mu$ with respect to $\nu$, such that each $\mu_{y}$ is a uniformly distributed probability measure on $F_{y}$. However, note that in this case the relation arc length and $\theta$ is no longer linear (see Figure 3).
Figure 3. Normalized arc length as a function of $\theta$ for $\lambda$ equals to $1,1.1,1.5$ and $2$.
Therefore, the measures $\mu_{y}$ have a different dependence on $\theta$:
$d\mu_{y}(r,\theta)=\frac{\frac{dL_{y}(\theta)}{d\theta}}{\mathcal{L}_{y}}~\delta_{y}~d\theta$
where
$L_{y}(\theta)=\int_{0}^{\theta}y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\tilde{\theta})}~d\tilde{\theta}$
is the arc length of the ellipse $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$ for angle between 0 and $\theta$, and $\mathcal{L}_{y}=L_{y}(2\pi)$. By the Fundamental Theorem of Calculus
$d\mu_{y}(r,\theta)=\frac{y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\theta)}}{\mathcal{L}_{y}}~\delta_{y}~d\theta.$
In this case, $d_{Y}(y,y^{\prime})=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\frac{1}{\lambda}|y-y^{\prime}|$ and, to calculate the $1$-Wasserstein distance between $\mu_{y}$ and $\mu_{y^{\prime}}$, we still have the radial transport, although $d(x,T(x))$ depends of $\theta$: for $x=y\cos(\theta)$, $d(x,T(x))=R_{y^{\prime}}(\theta)-R_{y}(\theta)$, where
$R_{y}(\theta)=\frac{y}{\sqrt{\lambda^{2}-(\lambda^{2}-1)\cos^{2}(\theta)}}$
and
$R_{y^{\prime}}(\theta)=\frac{y^{\prime}}{\sqrt{\lambda^{2}-(\lambda^{2}-1)\cos^{2}(\theta)}}.$
Then,
$W_{1}(\mu_{y},\mu_{y^{\prime}})=$ $\int R_{y^{\prime}}(\theta)-R_{y}(\theta)~d\mu_{y}$
$=$ $\int\frac{y^{\prime}-y}{\sqrt{\lambda^{2}-(\lambda^{2}-1)\cos^{2}(\theta)}}\frac{y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\theta)}}{\mathcal{L}_{y}}~\delta_{y}~d\theta$
$=$ $\frac{(y^{\prime}-y)~y}{\lambda~\mathcal{L}_{y}}~\int_{0}^{2\pi}~d\theta$
$=$ $\frac{2\pi~(y^{\prime}-y)~y}{\lambda~\mathcal{L}_{y}}.$
Therefore, considering $\hat{f}_{\lambda}$ the disintegration map of $\mu$ with respect to $\nu$, we have
$\mathcal{E}_{1}(\hat{f}_{\lambda})=\frac{2\pi y}{\mathcal{L}_{y}},$
that is, the energy is the ratio between the length of the circle with radius $y$ and the length of the ellipse with major axis equal to $y$. Table 1 displays the values of $\mathcal{L}_{y}$ for $\lambda=1$ (circle), $\lambda=1,001$, $\lambda=1,01$, $\lambda=1,1$, $\lambda=1,5$ and $\lambda=2$. Table 2 summarizes the computed values of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ corresponding to the cases in Table 1. Observe that $\mathcal{E}_{1}(\hat{f}_{\lambda})$ is independent of $y$.
Table 1. $\mathcal{L}_{y}$ for different values of $\lambda$
y $\lambda=1$ $\lambda=1,001$ $\lambda=1,01$ $\lambda=1,1$ $\lambda=1,5$ $\lambda=2$
11 $6,28318531$ $6,28004725$ $6,25211912$ $6,00098645$ $5,28847986$ $4,84422411$
Table 2. Values of $\mathcal{E}_{1}(\hat{f}_{\lambda})$
$\lambda=1$ $\lambda=1,001$ $\lambda=1,01$ $\lambda=1,1$ $\lambda=1,5$ $\lambda=2$
11 $1,000499687$ $1,004968906$ $1,047025412$ $1,188089106$ $1,297046785$
Figure 4 depicts the dependence of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ on $\lambda$. As shown, $\mathcal{E}_{1}$ grows smoothly with increasing foliation perturbation.
Figure 4. $\mathcal{E}_{1}(\hat{f}_{\lambda})$ as a function of $\lambda$
This example highlights the versatility of our approach: beyond classifying foliations on $X$ via disintegration maps, it also provides a framework for analyzing structural perturbations in foliations. Notably, the concepts developed here have potential applications across dynamical systems, machine learning, and stochastic analysis.
Acknowledgements
This work was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. C. S. R. has been partially supported by São Paulo Research Foundation (FAPESP): grant #2016/00332-1, grant #2018/13481-0, and grant #2020/04426-6. The opinions, hypotheses and conclusions or recommendations expressed in this work are the responsibility of the authors and do not necessarily reflect the views of FAPESP.
C. S. R. would like to acknowledge support from the Max Planck Society, Germany, through the award of a Max Planck Partner Group for Geometry and Probability in Dynamical Systems. R. P. would like to acknowledge support from the Max Planck Institute for Mathematics in the Sciences and from CAPES.

### v2

@@
-Classification of foliations via disintegration maps
+Characterization of foliations via disintegration maps
 Abstract.
@@
 Our method establishes criteria to determine when such conditional measures arise from a metric measure foliation.
-Additionally, we provide a concrete example demonstrating how this framework can be applied to study deformations (or perturbations) of disintegration-induced foliations.
+Additionally, we provide a example demonstrating how this framework can be applied to study perturbations of disintegration-induced foliations.
 1. Introduction
 The disintegration of measures offers a powerful framework for analyzing local and global probabilistic structures by decomposing a measure into conditional probabilities.
-Roughly speaking, given $X,Y$ measurable spaces, $\mu$ a measure on $X$, and a measurable map $\pi:X\to Y$, a disintegration allows us to break $\mu$ into a family of conditional measures $\{\mu_{y}\}_{y\in Y}$ on $X$, depending measurably on $y$, such that $\mu$ can be reconstructed by integrating $\mu_{y}$ against $\pi_{*}\mu$, the push-forward of $\mu$ by $\pi$.
+Roughly speaking, given measurable spaces $X,Y$, a measure $\mu$ on $X$, and a measurable map $\pi:X\to Y$, a disintegration allows us to break $\mu$ into a family of conditional measures $\{\mu_{y}\}$ on $X$, depending measurably on $y$, such that $\mu$ can be reconstructed by integrating $\mu_{y}$ against $\pi_{*}\mu$, the push-forward of $\mu$ by $\pi$.
 This decomposition is crucial in several areas, such as Probability Theory, Ergodic Theory, and Geometric Measure Theory, where it provides a rigorous way to handle conditional distributions.
-Disintegration theorems typically require topological assumptions to ensure existence and uniqueness of such decompositions, making them a powerful tool to define measures on fibered spaces or under projections, for instance.
+Disintegration theorems typically require topological assumptions to ensure existence and uniqueness of such decomposition, making them a key instrument to define measures on fibered spaces or under projections, for instance.
 Although significant progress has been made in establishing the existence of conditional measures under several different hypothesis, most of the existing approaches neglect the geometric aspects of the underlying space.
-In a recent paper [PR25], the second and third authors used the intrinsic structures of probability spaces to tackle the disintegration of measures from a perspective which takes into accounting statistical and geometrical properties of dynamical systems.
-In particular, it is introduced the notion of disintegration maps, which are essentially functions that assign to each point $y\in Y$ its corresponding conditional measure $\mu_{y}$, providing a structured way to analyze a disintegration.
-We emphasize that these objects can trace geodesics in the space of probability measures, which correspond to the solutions of a time-evolving transport problem, and exhibit certain rigidity phenomena: if one of the endpoints measure in the geodesic path is absolutely continuous, then so are all.
+In a recent paper [PR25], the second and third authors used the intrinsic structures of probability spaces to tackle the disintegration of measures from a perspective which takes into account statistical and geometrical properties of dynamical systems.
+In particular, it introduced the notion of disintegration maps, which are essentially functions that assign to each point $y\in Y$ its corresponding conditional measure $\mu_{y}$, providing a structured way to analyze a disintegration.
+We emphasize that these objects can trace geodesics in the space of probability measures, which correspond to the solutions of a time-evolving transport problem, and exhibit certain rigidity phenomena: if one of the endpoints measure in the geodesic path is absolutely continuous, then so are all inner points of the geodesic.
 Therefore, it becomes evident that certain geometric properties of the base space control the regularity of the conditional measures.
-In this paper, we further explore disintegration maps by defining a notion of derivative for these objects.
-Our notion of derivative resembles a metric derivative, comparing in a appropriated way the distance between conditional measures and the distance between the supports of such measures.
-We remark that, in order to define this notion of derivative we do not require any differentiable structure on the base space.
-An interesting case arises in the context of Riemannian submersions, for example, since the distance between fibers equals the distance between base points.
-In this context, if each conditional measure has full support on its corresponding fiber, we show that the distance among conditional measures and the distance between base points are the same, and this isometry is captured by the derivative.
-We introduce a novel dense energy functional, similarly to what is done in geometric analysis.
-This notion is suited to quantify the relationship between these distances more accurately.
-We show that the minimization of this functional detect and characterize highly ordered geometric structures in the base space.
-More specifically, we show that the minimum of this energy functional corresponds to a metric measure foliation.
-That is, a foliation of the base space whose leaves are parallel to each other and where the distance between conditional measures on these leaves equals the distance between the leaves themselves.
-Moreover, this functional exhibits sensitivity to small deformations of the foliation, capturing fine geometric structures.
-We provide an explicit example in which we assess the sensitivity of the energy functional by inducing deformations in a metric measure foliation.
-The main contributions of this paper are presented in Section 4, where we introduce a novel notion of derivative for disintegration maps (Definition 4.1).
-We then construct an energy functional designed to quantify the relationship between the Wasserstein distance of conditional measures and the distance between their supports (Definition 4.2).
+In this paper, we further explore disintegration maps by defining a notion of derivative for them.
+This derivative resembles a metric derivative, comparing the distance between conditional measures and the distance between the supports of such measures, without requiring a differentiable structure on the base space.
+Motivated by this derivative, we introduce a novel dense energy functional, similarly to what is done in Geometric Analysis, for cases in which the disintegration applies to every $y\in Y$.
+This functional characterizes highly ordered geometric structures in the base space: its minimum corresponds precisely to a metric measure foliation, that is, a foliation of the base space whose leaves are parallel to each other and where the distance between conditional measures on these leaves equals the distance between the leaves themselves.
+Metric measure foliation arises in various contexts, including Riemannian submersions between weighted manifolds, isometry group actions, and warped products, for instance [GKMS18, Kaz22].
+They are significant in Differential Geometry and Analysis due to their role in studying geometric properties, and providing a framework for convergence theory.
+One the one hand, metric foliations themselves represent a smooth partition of a Riemannian manifold into lower-dimensional submanifolds that are locally equidistant, similar to orbits of an isometric group action or parallel hypersurfaces [GG88].
+On the other hand, metric measure foliations are instrumental in the convergence theory of metric measure spaces, for instance.
+Research indicates that properties like the curvature-dimension condition and the Cheeger energy functional can be preserved from the original space to its quotient via these foliations.
+This framework also aids in investigating the convergence of sequences of metric measure spaces with unbounded dimensions [Kaz22].
+The main contributions of this paper are presented in Section 4, where we introduce a notion of derivative for disintegration maps (Definition 4.1) and an energy functional $\mathcal{E}_{p}$ (Definition 4.2), designed to quantify the relationship between the Wasserstein distance of conditional measures and the distance between their supports.
 Our main result is:
 Theorem 1.1.
-The minimum of $\mathcal{E}(f)$, the energy of the disintegration map, is equals to one, if and only if, the supports of conditional measures form a metric measure foliation.
+Given $f$ a disintegration map, the energy $\mathcal{E}_{p}(f)$ is equal to one precisely when the conditional measures have supports constituting a metric measure foliation.
 The precise statement shall be given in Theorem A.
 The paper is organized as follows.
-Section 2 outlines the theoretical framework and necessary definitions.
-Section 3 reviews essential background on disintegration maps.
-Finally, Section 4 presents our core results, including a detailed example exploring the sensitivity of the energy functional to deformations of the leaves that support the conditional measures.
+In Section 2, we establish the theoretical framework and requisite definitions.
+Section 3 reviews the essential background on disintegration maps.
+Our core results are presented in Section 4.
+These include examples that motivate the hypothesis in Theorem A, and a further example applying the energy functional to analyze the evolution of a measure under a flow.
 2. Theoretical framework
 Let $X$, $Y$ be locally compact complete separable metric spaces.
-Consider $\pi:X\to Y$ a Borel map, and $\mu\in\mathcal{M}_{+}(X)$, where $\mathcal{M}_{+}(X)$ is the set of all finite positive Radon measures on $X$.
+Consider a Borel map $\pi:X\to Y$, and $\mu\in\mathcal{M}_{+}(X)$, where $\mathcal{M}_{+}(X)$ is the set of all finite positive Radon measures on $X$.
 Define $\nu:=\pi_{*}\mu$ in $\mathcal{M}_{+}(Y)$.
-Then, by [PR25, Theorem A] there exist measures $\mu_{y}\in\mathcal{M}_{+}(X)$, such that,
-  1. (1)
-    $y\mapsto\mu_{y}$ is a Borel map for $\nu$-almost every $y\in Y$;
+Then, by [PR25, Theorem A], there exist measures $\mu_{y}\in\mathcal{M}_{+}(X)$ such that
   1. (1)
@@
   3. (3)
+    $y\mapsto\mu_{y}$ is a Borel map for $\nu$-almost every $y\in Y$; and
+  4. (4)
     $\mu(A)=\int_{Y}\mu_{y}(A)~d\nu(y)$ for every Borel subset $A$ of $X$.
-The family $\{\mu_{y}\}_{y\in Y}$ is called the disintegration of $\mu$ with respect to $\nu$.
-in order to examine this family of probability measures, we endow $\mathscr{P}(X)$, the space of probability measures on $X$, with a metric structure.
+The family $\{\mu_{y}\}$ is called the disintegration of $\mu$ with respect to $\nu$, and the measures $\mu_{y}$ are called conditional probabilities.
+In essence, this disintegration theorem provides a specific and constructive way to define conditional probabilities when the conditioning is based on a measurable function.
+This construction leads to more regular conditional probabilities, a property not universally guaranteed by the more abstract definition of conditional probability with respect to an arbitrary sub-$\sigma$-algebra.
+In order to examine this family of probability measures, we endow the space $\mathscr{P}(X)$ of probability measures on $X$, with a metric structure.
 The framework of Wasserstein spaces provides a natural geometric setting for such investigations.
 The Wasserstein space of order $p$, for $p\in[1,\infty)$, is the set
-$\mathscr{P}_{p}(X):=\Big{\{}\mu\in\mathscr{P}(X):\int d(x,\tilde{x})^{p}\mu(dx)<+\infty\Big{\}}$
-$\tilde{x}\in X$ arbitrary, endowed with the Wasserstein distance
-$W_{p}(\mu,\nu):=\Big{(}\inf\limits_{\gamma\in\Pi(\mu,\nu)}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})\Big{)}^{\frac{1}{p}}.$
-for $\mu$ and $\nu$ probability measures on $X$, where $\Pi(\mu,\nu)$ is the set of measures $\gamma\in\mathscr{P}(X\times X)$ with marginals $\mu$ and $\nu$.
-Given a disintegration of $\mu$ with respect to $\nu$, we call disintegration map a function assigns to each $y\in Y$ a conditional measure $\mu_{y}\in\mathscr{P}_{p}(X)$.
+$\mathscr{P}_{p}(X):=\Big\{\mu\in\mathscr{P}(X):\int d(x,\tilde{x})^{p}\mu(dx)<+\infty\Big\},$
+with $\tilde{x}\in X$ arbitrary, endowed with the Wasserstein distance
+$W_{p}(\mu,\nu):=\Big(\inf\limits_{\gamma\in\Pi(\mu,\nu)}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})\Big)^{\frac{1}{p}},$
+for probability measures $\mu$ and $\nu$ on $X$, where $\Pi(\mu,\nu)$ is the set of measures $\gamma\in\mathscr{P}(X\times X)$ with marginals $\mu$ and $\nu$.
+In this context, given a disintegration of $\mu$ with respect to $\nu$, by disintegration map we mean a function that assigns to each $y\in Y$ a conditional measure $\mu_{y}\in\mathscr{P}_{p}(X)$.
 Definition 2.1.
 [PR25, Definition 4.1] Let $X$, $Y$ be locally compact complete separable metric spaces.
-Consider $\pi:X\to Y$ a Borel map, $\mu\in\mathcal{M}_{+}(X)$, and $\nu:=\pi_{*}\mu$.
-Given $\{\mu_{y}\}_{y\in Y}$ a disintegration of $\mu$ with respect to $\nu$, we define the disintegration map by
-$f:D\subset Y$ $\to(\mathscr{P}_{p}(X),W_{p})$
+Consider a Borel map $\pi:X\to Y$, $\mu\in\mathcal{M}_{+}(X)$, and $\nu:=\pi_{*}\mu$.
+Given a disintegration $\{\mu_{y}\}$ of $\mu$ with respect to $\nu$, we define the disintegration map by
+$f:Y$ $\to(\mathscr{P}_{p}(X),W_{p})$
 $y$ $\mapsto\mu_{y},$
@@
 In order to make clear which measures are associated with the disintegration map, we say that “$f$ is a disintegration map of $\mu$ with respect to $\nu$”.
-To give an example, consider $X=[0,1]\times[0,1]$ the unit square, $Y=[0,1]$, $\mu=\text{Leb}^{2}$ the two-dimensional Lebesgue measure on $X$, $\pi(x,y)=x$ the projection onto the first coordinate, and $\nu=\text{Leb}^{1}$ the one-dimensional Lebesgue measure on $Y$.
-The disintegration $\{\mu_{x}\}_{x\in[0,1]}$ of $\mu$ with respect to $\nu$ is given by:
-$\mu_{x}=\text{Leb}^{1}\text{ on }\{x\}\times[0,1]\quad\text{(vertical line segment at }x\text{)}$
-Explicitly, for any measurable $A\subseteq X$:
-$\mu(A)=\int_{[0,1]}\mu_{x}(A)\,d\nu(x)=\int_{0}^{1}\text{Length}\big{(}A\cap(\{x\}\times[0,1])\big{)}\,dx$
-where $\text{Length}(A\cap(\{x\}\times[0,1])):=\text{Leb}^{1}(y\in[0,1]~:~(x,y)\in A)$.
-In fact, note that the map $x\mapsto\mu_{x}(A)=\text{Length}(A\cap\{x\}\times[0,1])$ is measurable, for any $B\subseteq[0,1]$,
-$\mu(A\cap\pi^{-1}(B))=\int_{B}\text{Length}\big{(}A\cap\{x\}\times[0,1]\big{)}\,dx$
-and $\mu_{x}$ is supported on $\pi^{-1}(x)=\{x\}\times[0,1]$, satisfying the disintegration conditions.
-Figure 1. Disintegration of $\text{Leb}^{2}$ with respect to $\text{Leb}^{1}$: the red line $\{x\}\times[0,1]$ carries $\mu_{x}=\text{Leb}^{1}$
+To give an example, consider the unit square $X=[0,1]\times[0,1]$, the two-dimensional Lebesgue measure $\mu=\text{Leb}^{2}$ on $X$, the projection onto the first coordinate $\pi(x,y)=x$, and the one-dimensional Lebesgue measure $\nu=\text{Leb}^{1}$ on $[0,1]$.
+A disintegration of $\mu$ with respect to $\nu$ is a family of measures $\mu_{x}=\text{Leb}^{1}\text{ on }\{x\}\times[0,1]$, with $x\in[0,1]$.
+Explicitly, for any measurable subset $A\subseteq X$,
+$\mu(A)=\int_{[0,1]}\mu_{x}(A)\,d\nu(x)=\int_{0}^{1}\text{Length}\big(A\cap(\{x\}\times[0,1])\big)\,dx,$
+where $\text{Length}(A\cap(\{x\}\times[0,1]))=\text{Leb}^{1}(\{y\in[0,1]:(x,y)\in A\})$.
+Moreover, the map $x\mapsto\mu_{x}(A)$ is measurable, and $\mu_{x}$ is supported on $\pi^{-1}(x)$, satisfying the disintegration conditions.
+Figure 1. Disintegration of $\text{Leb}^{2}$ with respect to $\text{Leb}^{1}$: the red line $\{x\}\times[0,1]$ carries $\mu_{x}=\text{Leb}^{1}$.
 In this case, the disintegration map is given by
@@
 $x$ $\mapsto\mu_{x},$
-where $\mu_{x}$ is $\text{Leb}^{1}$ on $\{x\}\times[0,1]$. This is a basic example that fits into the broader framework of metric foliations, which is the primary focus of our study.
-Let $(X,d)$ be a metric space, and $\mathcal{F}$ a partition of $X$ into closed subsets.
-We call $\mathcal{F}$ a foliation of $X$ and the elements of this partition are called leaves.
-If for every $F,F^{\prime}\in\mathcal{F}$ and every $x\in F$, $d(F,F^{\prime})=d(x,F^{\prime})$, where $d(F,F^{\prime})=\inf\{d(x,x^{\prime}):x\in F,x^{\prime}\in F^{\prime}\}$ and $d(x,F^{\prime})=d(\{x\},F^{\prime})$, then $\mathcal{F}$ is called a metric foliation.
+where $\mu_{x}$ is $\text{Leb}^{1}$ on $\{x\}\times[0,1]$.
+This is a basic example that fits into the broader framework of metric foliations, which is the primary focus of our study.
+Consider a metric space $(X,d)$, and a partition $\mathcal{F}$ of $X$ into closed subsets.
+We call $\mathcal{F}$ a foliation of $X$, and the elements of this partition are called leaves.
+If for every $\in\mathcal{F}, FF^{\prime}\in\mathcal{F}$, and every $x\in F$, holds $d(F,F^{\prime})=d(x,F^{\prime})$, where $d(F,F^{\prime})=\inf\{d(x,x^{\prime}):x\in F,x^{\prime}\in F^{\prime}\}$ and $d(x,F^{\prime})=d(\{x\},F^{\prime})$, then $\mathcal{F}$ is called a metric foliation.
 In case that each leaf is bounded, we say that $\mathcal{F}$ is bounded.
@@
 $x\sim x^{\prime}~\iff~\exists~F\in\mathcal{F}~\text{such that}~x,x^{\prime}\in F.$ (1)
-Consider the $X^{*}:=X/\sim$ the set of equivalence classes under (1) and the projection $\pi:X\to X^{*}$ onto $X^{*}$.
+Consider the set of equivalence classes under (1) $X^{*}:=X/\sim$ and the projection $\pi:X\to X^{*}$ onto $X^{*}$.
 We call $X^{*}$ the quotient space and $\pi$ the quotient map.
-Define a distance function $d^{*}$ on $X^{*}$ as
-$d^{*}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))$ (2)
-for $y,y^{\prime}\in X^{*}$. Note that $\pi$ is a submetry: $\pi(B(x,r))=B(\pi(x),r)$, where $B(x,r)$ is a ball centred at $x$ with radius $r$. In fact,
+We define $d^{*}$ a distance on $X^{*}$ by
+$d^{*}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime})).$ (2)
+Note that $\pi$ is a submetry: $\pi(B(x,r))=B(\pi(x),r)$, where $B(x,r)$ is a ball centered at $x$ with radius $r$. Indeed,
 $B(\pi(x),r)=$ $~\{y\in X^{*}:d^{*}(y,\pi(x))<r\}$
@@
 $=$ $~\pi(B(x,r)).$
-Therefore, $\pi$ is 1-Lipschitz. In this notation, we define a metric measure foliation.
+When the disintegration of $\mu$ with respect to a metric foliation $\mathcal{F}$ is such that the $p$-Wasserstein distance between two conditional measures $\mu_{y}$ and $\mu_{y^{\prime}}$ equals $d^{*}(y,y^{\prime})$, we say that $\mathcal{F}$ is a $p$-metric measure foliation.
+More precisely:
 Definition 2.2.
 Let $X$ be locally compact complete separable metric space, and $\mathcal{F}$ be a metric foliation of $X$.
-Consider $\mu\in\mathscr{M}_{+}(X)$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$.
-We call $\mathcal{F}$ a $p$-metric measure foliation, $p\in[1,\infty)$, if $\pi_{*}\mu$ is locally finite Borel measure on $X^{*}$, and
+Consider $\mu\in\mathscr{M}_{+}(X)$, the quotient space $X^{*}$ as defined above, and the quotient map $\pi:X\to X^{*}$.
+We call $\mathcal{F}$ a $p$-metric measure foliation, for $p\in[1,\infty)$, if $\pi_{*}\mu$ is locally finite Borel measure on $X^{*}$, and
 $W_{p}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})$ (3)
-for any $y,y^{\prime}\in Y$, where $\{\mu_{y}\}_{y\in Y}$ is a disintegration of $\mu$ with respect to $\pi_{*}\mu$.
+for any $y,y^{\prime}\in X^{*}$, where $\{\mu_{y}\}$ is a disintegration of $\mu$ with respect to $\pi_{*}\mu$.
 Remark 2.3.
-The general definition for a metric measure foliation requires $p=2$ and (3) only almost everywhere (see [GKMS18]). However, our definition is more restrictive, requiring (3) everywhere.
-Note that, in this case, the disintegration map is an isometry.
-This is the most important characteristic of metric measure foliation for us.
+The general definition for a metric measure foliation requires $p=2$ and (3) almost everywhere; see [GKMS18], for instance.
+However, our definition is more restrictive, requiring (3) everywhere.
+This restriction is incorporated into the definition in anticipation of the characterization established in Theorem A.
 Although the definition holds for $p\in[1,\infty)$, our case of interest is $p=2$, since $\mathscr{P}_{2}(X)$ has a rich geometric structure [GRS25].
-Moreover, if $\mathcal{F}$ is a 2-metric measure foliation, then
-$W_{q}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))$
-for any $q\in[1,\infty)$ and any $y,y^{\prime}\in X^{*}$, where $\{\mu_{y}\}_{y\in Y}$ is the disintegration of $\mu$ with respect to $\pi_{*}\mu$ [Kaz22, Proposition 3.11].
+Moreover, if $\mathcal{F}$ is a 2-metric measure foliation, then $W_{q}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})$, for any $q\in[1,\infty)$ and every $y,y^{\prime}\in X^{*}$, where $\{\mu_{y}\}$ is the disintegration of $\mu$ with respect to $\pi_{*}\mu$ [Kaz22, Proposition 3.11].
 Henceforth, the term “metric measure foliation” will refer to what we have defined as a “2-metric measure foliation”.
-A very important example of metric measure foliation is related to the action of isometry group [GR23]. Let $(X,d,\mu)$ be a metric measure space and $G$ a compact topological group. Let
+A very important example of metric measure foliation is related to the action of isometry group [GR23]. Let $(X,d,\mu)$ be a metric measure space and consider a compact topological group $G$. Let
 $G\times X\ni(g,x)\mapsto gx\in X$
@@
 That is, for every $g\in G$ the map $X\ni x\mapsto gx\in X$ is an isometry preserving the measure $\mu$.
-Consider $[x]$ the $G$-orbit of a point $x\in X$ and the quotient space $X/G$ endowed with the distance
+Consider the $G$-orbit $[x]$ of a point $x\in X$ and the quotient space $X/G$ endowed with the distance
 $d_{X/G}([x],[x^{\prime}])=\inf_{g,g^{\prime}\in G}d(gx,g^{\prime}x^{\prime}).$
-Consider $\pi:X\to X/G$ the projection map, that is, $\pi$ is given by $x\mapsto[x]$. The family $\mathcal{F}:=\{\pi^{-1}(y):y\in X/G\}$ is a metric measure foliation on $X$.
-Other interesting examples arise from Riemannian submersions of weighted Riemannian manifolds [GKMS18], for instance.
+Consider the projection map $\pi:X\to X/G$, that is, $\pi$ is given by $x\mapsto[x]$.
+The family $\mathcal{F}:=\{\pi^{-1}(y):y\in X/G\}$ is a metric measure foliation on $X$.
+Other interesting examples arise from Riemannian submersions of weighted Riemannian manifolds; see [GKMS18], for instance.
 3. Preliminaries on disintegration maps
 In this section, we discuss key properties of disintegration maps, as introduced in Definition 2.1 and developed in [PR25].
-To start our discussion, suppose $X=Y=\mathbb{R}$, and let $\gamma$ be a probability measure supported on the graph of a measurable function $\pi:X\to Y$, i.e., $\gamma=(\mathrm{id},\pi)_{*}\mu$, where $\mu$ its a probability measure on $X$ and $(\mathrm{id},\pi)_{*}\mu$ is the pushforward of $\mu$ under the map $x\mapsto(x,\pi(x))$.
-Then, the disintegration of $\gamma$ with respect to $\mu$ is given by the family of Dirac measures:
-$\gamma_{x}=\delta_{\pi(x)}.$
-In fact, for any measurable $A\subseteq X\times Y$, we have
+To start our discussion, suppose $X=Y=\mathbb{R}$, consider a measurable map $\pi:X\to Y$ and let $\gamma$ be a probability measure supported on the graph of $\pi$, i.e., $\gamma=(\mathrm{id},\pi)_{*}\mu$, where $\mu$ is some probability measure on $X$ and $(\mathrm{id},\pi)_{*}\mu$ is the push-forward of $\mu$ under the map $x\mapsto(x,\pi(x))$.
+Then, the disintegration of $\gamma$ with respect to $\mu$ is given by a family of Dirac measures $\gamma_{x}=\delta_{\pi(x)}$.
+Indeed, for every measurable set $A\subseteq X\times Y$, one has
 $\gamma(A)=\mu(\{x\in X:(x,\pi(x))\in A\}).$
-By definition of the disintegration
-$\gamma(A)=\int_{X}\gamma_{x}(A_{x})\,d\mu(x)=\int_{X}\delta_{\pi(x)}(A_{x})\,d\mu(x).$
-where $A_{x}=\{y\in Y:(x,y)\in A\}$. Then,
+By definition of the disintegration, it holds
+$\gamma(A)=\int_{X}\gamma_{x}(A_{x})\,d\mu(x)=\int_{X}\delta_{\pi(x)}(A_{x})\,d\mu(x),$
+with $A_{x}=\{y\in Y:(x,y)\in A\}$. Then,
 $\delta_{\pi(x)}(A_{x})=\mathbbm{1}_{A}(x,\pi(x)),$
@@
 Observe that when $\pi=\mathrm{id}$, we have a natural isometric immersion of $X$ into $\mathscr{P}_{p}(X)$, namely the disintegration map $x\mapsto\delta_{x}$.
-In general, we aim to investigate under what conditions geometric information about $X$ can be derived via disintegration — specifically, through the properties of disintegration maps.
+In general, we aim to investigate under what conditions geometric information about $X$ can be derived via disintegration, specifically, through the properties of disintegration maps.
 A key starting point for this analysis is understanding the continuity of these maps.
@@
     [PR25, Propostion 5.2] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$.
-If $\pi:X\to Y$ is a Borel map such that $\nu:=\pi_{*}\mu$ is a Borel measure, then, the disintegration map $f:Y\to(\mathscr{P}_{p}(X),W_{p})$ of $\mu$ with respect to $\nu$ is nearly continuous, for $p\in[1,\infty)$.
-A map $f$ on an metric space $(Y,\nu)$ is called nearly continuous if, for each $\varepsilon>0$ there exist $\mathcal{K}\subset Y$ closed with $\nu(Y\backslash\mathcal{K})<\varepsilon$ such that $f$ restricted to $\mathcal{K}$ is continuous.
+If a Borel map $\pi:X\to Y$ is such that $\nu:=\pi_{*}\mu$ is a Borel measure, then, the disintegration map $f:Y\to(\mathscr{P}_{p}(X),W_{p})$ of $\mu$ with respect to $\nu$ is nearly continuous, for $p\in[1,\infty)$.
+Here, a map $f$ on an metric space $(Y,\nu)$ is called nearly continuous if for each $\varepsilon>0$ there exist a closed subset $\mathcal{K}\subset Y$ with $\nu(Y\backslash\mathcal{K})<\varepsilon$ such that $f$ restricted to $\mathcal{K}$ is continuous.
   2. (2)
     [PR25, Propostion 5.4] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$.
-If the Borel map $\pi:X\to Y$ is bijective and $\pi^{-1}$ continuous, then the disintegration map $f$ of $\mu$ with respect to $\pi_{*}\mu$ is continuous.
+If a Borel map $\pi:X\to Y$ is bijective and $\pi^{-1}$ is continuous, then the disintegration map $f$ of $\mu$ with respect to $\pi_{*}\mu$ is continuous.
   3. (3)
-    [PR25, Propostion 5.8] Let $X$ be locally compact complete separable metric space.
-Considering a metric foliation $\mathcal{F}$ of $X$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$, if there exists a metric measure foliation of $X$, then the disintegration map of $\mu$ with respect to $\nu:=\pi_{*}\mu$ is continuous.
+    [PR25, Propostion 5.8] Let $X$ be a locally compact complete separable metric space.
+Consider a metric foliation $\mathcal{F}$ of $X$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$.
+If $\mathcal{F}$ is a metric measure foliation, then the disintegration map of $\mu$ with respect to $\nu:=\pi_{*}\mu$ is continuous.
 The continuity of disintegration maps enables us to relate paths in the chosen indexing metric space to corresponding paths of measures in Wasserstein space.
@@
 In the context of a metric measure foliation, this relationship becomes more nuanced.
-Here, transversal paths (those crossing the leaves) must satisfy an additional condition concerning the distance between the supports of the conditional measures along the path — a consequence of the parallelism of the leaves.
-In the next section we present an approach that allows us to study the relation between the relative position of the elements of the family of supports and the relative position of the conditional measures in the Wasserstein space, via a disintegration map.
+Here, transversal paths (those crossing the leaves) must satisfy an additional condition concerning the distance between the supports of the conditional measures along the path, a consequence of the parallelism of the leaves.
+In the next section we present an approach that allows us to study the relation between the relative position of the conditional measures in the Wasserstein space and the relative position of their respective supports, via disintegration map.
 4. On foliations and disintegration maps
-Motivated by the concept of the metric derivative and the interplay between the relative positions of supports and conditional measures, we introduce a notion of derivative for disintegration maps to quantify how these measures evolve along paths.
+In this section, we will assume that the disintegration of measures holds for every point, rather than for $\nu$-almost every point, denoting the family as $\{\mu_{y}\}_{y\in Y}$ to explicitly indicate that the conditional measures are defined for every $y\in Y$.
+In fact, under some structural assumptions, the disintegration can hold for every $y\in Y$.
+This may occurs, for example, when $\pi$ is a trivial bundle with compact fiber, or when $Y$ is discrete.
+In such settings, $\mu_{y}$ can be defined point wise while preserving all disintegration properties.
+Motivated by the concept of the metric derivative and the interplay between the relative positions of conditional measures and their supports, we introduce a notion of derivative for disintegration maps.
 Definition 4.1.
-Let $X,Y$ be locally compact complete separable metric spaces.
-Consider a measure $\mu\in\mathcal{M}_{+}(X)$, $\pi:X\to Y$ a Borel map, $\nu=\pi_{*}\mu$, and $f$ the disintegration map of $\mu$ with respect to $\nu$.
-We define a notion of derivative of $f$ by
-$|\nabla f(y)|_{p}:=\lim\limits_{\varepsilon\to 0}\sup_{y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)}\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))}$
-for $p\in[1,\infty)$, where $\mu_{y^{\prime}}=f(y^{\prime})$ and $\mu_{y^{\prime\prime}}=f(y^{\prime\prime})$ are conditional measures and $B_{\varepsilon}(y)$ is a ball centered in $y\in Y$ with radius $\varepsilon$.
-Let $d_{Y}(y^{\prime},y^{\prime\prime})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$ denote the distance between the fibers over $y^{\prime}$ and $y^{\prime\prime}$.
-In the case of a metric measure foliation, the $p$-Wasserstein distance between conditional measures satisfies $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d_{Y}(y^{\prime},y^{\prime\prime})$, which implies $|\nabla f(y)|_{p}=1$.
-This derivative concept thus serves as a tool to characterize when $\{\pi^{-1}(y)\}$ forms a metric measure foliation.
-To further analyze the connection between the distances of conditional measure supports and their Wasserstein distance, paralleling techniques in geometric analysis, we introduce a dense energy functional designed to quantify this relationship.
+Let $X$ and $Y$ be locally compact complete separable metric spaces.
+Consider a measure $\mu\in\mathcal{M}_{+}(X)$ and a Borel map $\pi:X\to Y$ such that $\{\mu_{y}\}$ is a disintegration of $\mu$ with respect to $\nu=\pi_{*}\mu$.
+Let $f$ be the respective disintegration map.
+We write, for $y^{\prime},y^{\prime\prime}\in Y$,
+$\rho(y^{\prime},y^{\prime\prime}):=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime})),$
+and we define a notion of derivative of $f$ by
+$|\nabla f(y)|_{p}:=\lim\limits_{\varepsilon\to 0}~\sup_{\begin{subarray}{c}\rho(y,y^{\prime})\leq\varepsilon\\ \rho(y,y^{\prime\prime})\leq\varepsilon\\ y^{\prime}\neq y^{\prime\prime}\end{subarray}}~\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{\rho(y^{\prime},y^{\prime\prime})}$ (4)
+for $p\in[1,\infty)$, where $\mu_{y^{\prime}}=f(y^{\prime})$ and $\mu_{y^{\prime\prime}}=f(y^{\prime\prime})$ are conditional measures.
+We warn the reader that $\rho$ may not satisfy all the distance axioms.
+Note that $|\nabla f(y)|_{p}\geq 1$ everywhere.
+Indeed, since $\mu_{y^{\prime}}$ and $\mu_{y^{\prime\prime}}$ are supported on $\pi^{-1}(y^{\prime})$ and $\pi^{-1}(y^{\prime\prime})$, respectively, we have
+$W_{p}^{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}}):=$ $\inf\limits_{\gamma\in\Pi(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})$
+$\geq$ $~d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))^{p}.$
+In the case of a $p$-metric measure foliation, the $p$-Wasserstein distance between conditional measures satisfies $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$, which implies $|\nabla f(y)|_{p}=1$ for every $y\in Y$.
+Meanwhile, $|\nabla f(y)|_{p}=1$ for every $y\in Y$ does not necessarily imply that the family of fibers $\{\pi^{-1}(y)\}$ is a $p$-metric measure foliation.
+However, this derivative can be used as a tool to characterize when $\{\pi^{-1}(y)\}$ forms a $p$-metric measure foliation, provided certain additional assumptions are made.
+Paralleling techniques from Geometric Analysis and Optimal Transport, we introduce a dense energy functional designed to further analyze the relationship between the distances of conditional measures and their supports.
 Definition 4.2.
-Let $X,Y$ be locally compact complete separable metric spaces.
-Consider a measure $\mu\in\mathcal{M}_{+}(X)$, a Borel map $\pi:X\to Y$, and the disintegration map $f$ of $\mu$ with respect to $\nu=\pi_{*}\mu$.
+Let $X$ and $Y$ be locally compact complete separable metric spaces.
+Consider a measure $\mu\in\mathcal{M}_{+}(X)$ and a Borel map $\pi:X\to Y$ such that $\{\mu_{y}\}_{y\in Y}$ is a disintegration of $\mu$ with respect to $\nu=\pi_{*}\mu$.
+Let $f$ be the respective disintegration map.
 The $p$-energy of $f$ is given by
 $\mathcal{E}_{p}(f):=\|\nabla f\|_{\infty,p}=\sup_{y\in Y}|\nabla f(y)|_{p}.$
-Note that $|\nabla f|_{p}\geq 1$ everywhere.
-In fact, since $\mu_{y^{\prime}}$ and $\mu_{y^{\prime\prime}}$ are supported on $\pi^{-1}(y^{\prime})$ and $\pi^{-1}(y^{\prime\prime})$, respectively, we have
-$W_{p}^{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}}):=$ $\inf\limits_{\gamma\in\Pi(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})$
-$\geq$ $~d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))^{p}.$
-Moreover, note that $\min_{f}\mathcal{E}_{p}(f)=1$ if, and only if, $\|\nabla f\|_{\infty,p}=1$, which in turn hold true if, and only if $|\nabla f|_{p}=1$.
-In this case we have an isometry $W_{p}=d_{Y}$.
-Under some further hypotheses, we can relate the functional $\mathcal{E}_{p}(f)$ to the metric measure foliation case.
+This energy definition involves analyzing $|\nabla f(y)|_{p}$ at every point $y\in Y$.
+Example 4.4 demonstrates a scenario in which the definition is satisfied for almost every point, yet the characterization we seek to prove does not hold.
+In the following theorem, we relate the value of this energy functional to settings in which the disintegration corresponds to a metric measure foliation case.
 Theorem A.
-Let $(X,d)$ be a locally compact complete separable geodesic space, $Y$ a compact complete separable metric space, and $\pi:X\to Y$ a Borel map.
-Given $\mu\in\mathcal{M}_{+}(X)$, let $f_{\mu,\pi}$ be a disintegration of $\mu$ with respect to $\nu:=\pi_{*}\mu$.
-Suppose that the conditional measures $\mu_{y}$ have full support, i.e., $\operatorname{supp}(\mu_{y})=\pi^{-1}(y)$.
-Then, $\mathcal{E}_{p}(f_{\mu,\pi})=1$ if, and only if, $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$.
-Therefore, the energy-minimizing disintegration maps correspond to a metric measure foliation.
+Let $(X,d_{X})$ be a geodesic locally compact complete separable metric space.
+Consider a locally compact complete separable metric space $Y$, a measure $\mu\in\mathcal{M}_{+}(X)$ and a Borel map $\pi:X\to Y$ such that $\{\mu_{y}\}_{y\in Y}$ is a disintegration of $\mu$ with respect to $\nu:=\pi_{*}\mu$, with $\operatorname{supp}(\mu_{y})=\pi^{-1}(y)$.
+Let $f$ be the respective disintegration map.
+Then, $\mathcal{E}_{p}(f)=1$ if and only if $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$.
+Remark 4.3.
+The assumptions on the space $Y$ are made only to guarantee the existence of the disintegration of $\mu$ with respect to $\pi_{*}\mu$, according [PR25, Theorem A].
+Indeed, the claim of our theorem only concerns the distance between the fibers, and not the original distance on $Y$.
 Proof.
-[…]
-$u(y^{\prime},y^{\prime\prime}):=\begin{cases}\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{d_{Y}(y^{\prime},y^{\prime\prime})}&~\text{,if}~~y^{\prime}\neq y^{\prime\prime}\\ 1&~\text{,if}~~y^{\prime}=y^{\prime\prime}\end{cases}$
-Note that $u\geq 1$. By the hypothesis,
-$\lim\limits_{\varepsilon\to 0}\sup\limits_{y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)}u(y^{\prime},y^{\prime\prime})=u(y,y)$
-for every $y\in Y$.
-Then, $u$ is continuous at $(y,y)$ for every $y\in Y$.
-Moreover, $u$ is continuous everywhere else.
-In fact, since $\min_{f_{\mu,\pi}}\mathcal{E}_{p}(f_{\mu,\pi})=1$, we have $|\nabla f(y)|p=1$ for every $y\in Y$, and the disintegration map $f$ is continuous.
-Thus, it implies the continuity of $u$ everywhere else.
-Since $Y$ compact, $Y\times Y$ is compact and by compactness, we have uniform continuity, that is, for all $\varepsilon>0$ there exists $\delta>0$ such that,
-$d_{Y\times Y}((y,y^{\prime}),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$
-implies $|u(y,y^{\prime})-u(\tilde{y},\tilde{y}^{\prime})|\leq\varepsilon$, for every ($y~)(y,y^{\prime}$),($\tilde$$\tilde{y}^{\prime}$), where
-[…]
-Note that $|u(y,y)-u(\tilde{y},\tilde{y}^{\prime})|=u(\tilde{y},\tilde{y}^{\prime})-1$.
-Then, setting $y=\tilde{y}$, and if $d_{Y}((\tilde{y},\tilde{y}^{\prime}),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$, we get $d_{Y\times Y}((y,y),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$, Thus,
-$|u(y,y)-u(\tilde{y},\tilde{y}^{\prime})|=u(\tilde{y},\tilde{y}^{\prime})-1\leq\varepsilon,$
-[…]
-$\sum_{k=0}^{n-1}d(x_{k},x_{k+1})=d(x,x^{\prime}).$
-Then,
-[…]
-$\leq$ $\sum_{k=0}^{n-1}d(\pi^{-1}(\pi(x_{k})),\pi^{-1}(\pi(x_{k+1})))$
-$\leq$ $(1+\varepsilon)\sum_{k=0}^{n-1}d(\pi^{-1}(\pi(x_{k})),\pi^{-1}(\pi(x_{k+1})))$
-$\leq$ $(1+\varepsilon)\sum_{k=0}^{n-1}d(x_{k},x_{k+1})$
-$=$ $(1+\varepsilon)d(x,x^{\prime})$
-[…]
-[…]
-[…]
-[…]
-Let $\gamma$ be the optimal transport plan from $\mu_{y}$ to $\mu_{y^{\prime}}$, and consider, for all $(x,x^{\prime})\in\operatorname{supp}(\gamma)$, $\varphi(x)=d(x,\pi^{-1}(y^{\prime}))>d(x,x^{\prime})$.
-Then,
-$\int\varphi(x)d\gamma=\int\varphi(x)d\mu_{y}=\int d(x,\pi^{-1}(y^{\prime}))d\mu_{y}>\int d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))d\mu_{y}$
-that is a contradiction, since
-$W_{1}(\mu_{\tilde{y}},\mu_{\tilde{y}^{\prime}}):=\inf\limits_{\gamma\in\Pi(\mu_{\tilde{y}},\mu_{\tilde{y}^{\prime}})}\int d(x_{1},x_{2})~d\gamma(x_{1},x_{2}).$
-An analogous argument applies to the case $p\in(1,\infty)$.
-Remark 4.3.
-The conclusion of Theorem A does not hold if full support is not assumed.
-In fact, consider $X=[-1,1]\times\mathbb{R}$, $Y=[0,1]$, $\pi:X\to Y$ such that $\pi^{-1}(y)$ is a ellipse given by $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$, with major radius $y$.
+The demonstration will be carried out in three steps.
+Step 1 shows that $p$-metric measure foliation implies $\mathcal{E}_{p}(f)=1$, and Step 2 and Step 3 show that if $\mathcal{E}_{p}(f)=1$, then $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$.
+More specifically, Step 2 shows that the Wasserstein distance between two conditional measures coincides with the distance between the corresponding fibers; Step 3 shows the metric foliation property, that is, all points on a fiber have the same distance to another fiber.
+In the sequel, we consider the distance between the preimages of $y^{\prime}$ and $y^{\prime\prime}$ to be given by $\rho(y^{\prime},y^{\prime\prime})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))=\inf\{d_{X}(x^{\prime},x^{\prime\prime}):\mbox{$x^{\prime}\in\pi^{-1}(y^{\prime})$}$, $x^{\prime\prime}\in\pi^{-1}(y^{\prime\prime})\}$.
+Step 1.
+Suppose that $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$.
+Then, for any $y^{\prime},y^{\prime\prime}\in Y$, holds $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$.
+Therefore, for any $y\in Y$, one has
+$|\nabla f(y)|_{p}=\lim\limits_{\varepsilon\to 0}~\sup_{\begin{subarray}{c}\rho(y,y^{\prime})\leq\varepsilon\\ \rho(y,y^{\prime\prime})\leq\varepsilon\\ y^{\prime}\neq y^{\prime\prime}\end{subarray}}~\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{\rho(y^{\prime},y^{\prime\prime})}=1,$
+which implies $\mathcal{E}_{p}(f)=1$.
+Step 2.
+Suppose $\mathcal{E}_{p}(f)=1$.
+We want to show that in this case, for any $y^{\prime},y^{\prime\prime}\in Y$, it holds that $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$.
+Consider $\varepsilon>0$ and $y_{0},y_{1}\in Y$.
+Then there exist $x_{0}\in\pi^{-1}(y_{0})$, $x_{1}\in\pi^{-1}(y_{1})$, such that $d_{X}(x_{0},x_{1})\leq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))+\varepsilon$.
+Let $\gamma:[0,1]\to X$ be a constant speed minimizing geodesic such that $\gamma(0)=x_{0}$ and $\gamma(1)=x_{1}$.
+Define $y_{t}:=\pi(\gamma(t))$, $t\in[0,1]$.
+Note that, for every $s,t\in[0,1]$, it holds
+$W_{p}(\mu_{y_{s}},\mu_{y_{t}})\geq d(\pi^{-1}(y_{s}),\pi^{-1}(y_{t}))$ (5)
+and
+$d(\pi^{-1}(y_{s}),\pi^{-1}(y_{t}))\leq|s-t|d_{X}(x_{0},x_{1}).$ (6)
+Define $\varphi(t):=W_{p}(\mu_{y_{t}},\mu_{y_{0}})$. We will show that $\varphi$ is continuous and that
+$\partial_{t}^{+}(\varphi(t)):=\limsup_{h\to 0}\frac{\varphi(t+h)-\varphi(t)}{h}\leq d_{X}(x_{0},x_{1}).$
+Indeed,
+$\limsup_{h\to 0}\left|\frac{\varphi(t+h)-\varphi(t)}{h}\right|=$ $\limsup_{h\to 0}\left|\frac{W_{p}(\mu_{y_{t+h}},\mu_{y_{0}})-W_{p}(\mu_{y_{t}},\mu_{y_{0}})}{h}\right|$
+$\leq$ $\limsup_{h\to 0}\frac{W_{p}(\mu_{y_{t+h}},\mu_{y_{t}})}{|h|}$
+$\leq$ $d_{X}(x_{0},x_{1})\limsup_{h\to 0}\frac{W_{p}(\mu_{y_{t+h}},\mu_{y_{t}})}{d(\pi^{-1}(y_{t+h}),\pi^{-1}(y_{t}))}$
+$\leq$ $d_{X}(x_{0},x_{1})|\nabla f(y_{t})|_{p}$
+$\leq$ $d_{X}(x_{0},x_{1})\mathcal{E}_{p}(f)$
+$=$ $d_{X}(x_{0},x_{1}).$
+Here, the first inequality follows from the triangle inequality for the Wasserstein distance; the second one follows from (6); the third one follows from the definition of the gradient (4); the last equation follows from the assumption $\mathcal{E}_{p}(f)=1$.
+Now, the continuity of $\varphi$ follows easily.
+By [HT06, Theorem 2], we can integrate the inequality $\partial_{t}^{+}(\varphi(t))\leq d_{X}(x_{0},x_{1})$, giving $\varphi(1)-\varphi(0)\leq d_{X}(x_{0},x_{1})$. Thus,
+$W_{p}(\mu_{y_{0}},\mu_{y_{1}})\leq d_{X}(x_{0},x_{1})\leq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))+\varepsilon.$
+Since $\varepsilon$ is arbitrary, taking the limit $\varepsilon\to 0$, we have
+$W_{p}(\mu_{y_{0}},\mu_{y_{1}})\leq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1})).$ (7)
+By (5) and (7) we obtain $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$.
+Step 3.
+Keep assuming $\mathcal{E}_{p}(f)=1$.
+We will show that for every $y_{0},y_{1}\in Y$, and for every $x_{0}\in\pi^{-1}(y_{0})$, it holds $d(x_{0},\pi^{-1}(y_{1}))=d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))$.
+Note that, by definition, one has $d(x_{0},\pi^{-1}(y_{1}))\geq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))$.
+Consider a function $r:\pi^{-1}(y_{0})\to\mathbb{R}^{+}_{0}$ given by $r(x):=d(x,\pi^{-1}(y_{1}))$.
+Note that $r(x)\geq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))$ for every $x\in\pi^{-1}(y_{0})$ and $r$ is continuous.
+Moreover, note that for $(x,x^{\prime})\in\pi^{-1}(y_{0})\times\pi^{-1}(y_{1})$, one has $d_{X}(x,x^{\prime})\geq r(x)$.
+Let $\eta$ be an optimal transport plan from $\mu_{y_{0}}$ to $\mu_{y_{1}}$.
+Then, by Step 2, it holds
+$\int d_{X}(x,x^{\prime})^{p}d\eta=W^{p}_{p}(\mu_{y_{0}},\mu_{y_{1}})=[d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))]^{p}.$
+Furthermore, one has
+$\int d_{X}(x,x^{\prime})^{p}d\eta\geq\int r(x)^{p}d\mu_{y_{0}}\geq[d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))]^{p}.$
+Since $r^{p}$ is continuous and $\operatorname{supp}(\mu_{y_{0}})=\pi^{-1}(y_{0})$, for every $x\in\pi^{-1}(y_{0})$ we have $r(x)^{p}=[d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))]^{p}$ .
+Therefore, for every $x_{0}\in\pi^{-1}(y_{0})$, we have $d(x_{0},\pi^{-1}(y_{1}))=d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))$
+Step 2 and Step 3 complete the proof.
+In Theorem A, we characterize a foliation on $X$ via the energy of a disintegration map.
+We emphasize that the hypotheses of conditional measures having full support and the analysis of the energy functional over all points are fundamental to the result.
+In the following examples, we present cases that justify the use of such assumptions.
+Example 4.4.
+We will show that there exists a disintegration $\{\mu_{y}\}$, such that $|\nabla f(y)|_{p}=1$ for $\nu$-almost every $y$, but the metric measure foliation condition does not hold.
+Consider X=[2,1]X=[-2,1], $Y=[0,1]$ and $\pi:X\to Y$ such that
+$\pi(x)=\begin{cases}x&,\text{if}~x\geq 0\\ (\phi+\text{id})^{-1}(-x)&,\text{if}~x<0\end{cases}$
+where $\phi$ is the Cantor ternary function.
+Consider a measure $\mu$ on $X$ such that $\nu=\pi_{*}\mu$ is a uniform measure on $Y$ and there exists a disintegration of $\mu$ with respect to $\nu$ such that the conditional measures are
+$\mu_{y}=\frac{1}{2}\delta_{y}+\frac{1}{2}\delta_{\varphi(y)}$
+where $\varphi(y)=\phi(y)+y$. Note that, for $y^{\prime},y^{\prime\prime}\in Y$,
+$d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))=$ $d(\{y^{\prime},\varphi(y^{\prime})\},\{y^{\prime\prime},\varphi(y^{\prime\prime})\})$
+$=$ $\min\left\{|y^{\prime}-y^{\prime\prime}|,|\varphi(y^{\prime}),\varphi(y^{\prime})|\right\}$
+$=$ $|y^{\prime}-y^{\prime\prime}|.$
+Moreover, let $\mathcal{U}$ be the union of the maximum intervals where the Cantor function is constant, i.e., $\mathcal{U}=(\frac{1}{3},\frac{2}{3})\cup(\frac{1}{9},\frac{2}{9})\cup(\frac{7}{9},\frac{8}{9})\cup\dots$.
+Note that $\mathcal{U}$ is dense in $[0,1]$, $\nu(\mathcal{U})=1$, and for $y\in\mathcal{U}$, for all $y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)$, $\varepsilon<d_{Y}(y,\mathcal{U}^{c})$, we have $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=|y^{\prime}-y^{\prime\prime}|$.
+Therefore, $|\nabla f(y)|_{p}=1$ for $\nu$-almost every $y\in Y$.
+Example 4.4 justifies the hypothesis of taking the supremum over all $y\in Y$ in Definition 4.2.
+If we had defined the energy functional using an essential supremum, for example, pathological cases like this would arise, where the other hypotheses of Theorem A are satisfied, but we do not have a $p$-metric measure foliation.
+Another important remark: when considering a general disintegration $\{\mu_{y}\}$, which is defined $\nu$-almost everywhere, we could define measures at the missing points $y$, but this does not alter the geometric structure on $X$.
+Moreover, note that by requiring that the conditional measures have full support, the fiber structure becomes detectable by the energy functional.
+The following example illustrates a case in which Theorem A does not hold if the full support condition is not assumed.
+Example 4.5.
+Consider $X=[-1,1]\times\mathbb{R}$, $Y=[0,1]$, $\pi:X\to Y$ such that $\pi^{-1}(y)$ is an ellipse given by $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$, $\lambda>1$, with major radius $y$.
 Consider a measure $\mu\in\mathcal{M}_{+}(X)$ such that the disintegration of $\mu$ with respect to $\nu$ is $\mu_{y}=\delta_{(0,\frac{y}{\lambda})}$.
 In this case,
-$d_{Y}(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\left|\frac{y}{\lambda}-\frac{y^{\prime}}{\lambda}\right|=W_{p}(\mu_{y},\mu_{y^{\prime}})$
-for every $y,y^{\prime}\in Y$. Therefore, we have isometry, but $\{\pi^{-1}(y)\}$ does not define a metric foliation. A graphical representation of the case $\lambda=1.5.$ is provided in Figure 2.
+$d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\left|\frac{y}{\lambda}-\frac{y^{\prime}}{\lambda}\right|=W_{p}(\mu_{y},\mu_{y^{\prime}})$
+for every $y,y^{\prime}\in Y$, but $\{\pi^{-1}(y)\}$ does not define a metric foliation. A representation of the case $\lambda=1.5.$ is provided in Figure 2.
 Figure 2.
-Illustration for the case in which the measures $\mu_{y}$ do not have full support: there is isometry between $W_{p}$ and $d_{Y}$, but $\{\pi^{-1}(y)\}$ (the ellipses) does not define a metric foliation.
-Note that in the previous example, if $\lambda=1$, that is, the preimages are circles in $X$, then $\{\pi^{-1}(y)\}$ forms a metric foliation.
-In the context of Theorem A, we aim to demonstrate that the $p$-energy exhibits sensitivity to small perturbations of the foliation.
-To analyze this, we first evaluate the $1$-energy $\mathcal{E}_{1}$ for a metric measure foliation on $\mathbbm{R}^{2}$ with concentric circular leaves.
-We then generalize this analysis to cases where the leaves are slightly deformed into ellipses.
+Illustration for the case in which the conditional measures do not have full support: $d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})$ for every $y^{\prime},y^{\prime\prime}\in Y$, but $\{\pi^{-1}(y)\}$ (the ellipses) are not a metric foliation.
+In the next example, we analyze the disintegration of a measure $\mu$ under a flow $\varphi$.
+For $t=0$, we have the characterization of a metric measure foliation given by Theorem A.
+We use the energy functional to analyze the change in the foliation associated with the measure $\mu_{t}$, which is the push-forward of $\mu$ by the flow $\varphi$ at time $t$.
+This example shows that the energy functional exhibits a certain sensitivity to smooth changes in the foliation structure.
 Example 4.6.
-Consider $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}\leq R^{2}\}$, $R>0$, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed, and $Y=(0,R]$.
+Consider $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}\leq R^{2}\}$, $R>0$, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed on $X$, and $Y=(0,R]$.
 Consider a foliation of $X$ into circles, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+x_{2}^{2}}=y$.
-To simplify the analysis, consider this setup in polar coordinates: consider $x_{1}=r\cos(\theta)$, $x_{2}=r\sin(\theta)$.
-Using Cartesian coordinates, an infinitesimal area element can be calculated as
-$dA=dx_{1}~dx_{2}.$
-To convert it to polar coordinates, the Jacobian determinant $J$ of the coordinate conversion formula has to be considered. In this case:
-$J(r,\theta)=\begin{vmatrix}\frac{\partial x_{1}}{\partial r}&\frac{\partial x_{1}}{\partial\theta}\\ \frac{\partial x_{2}}{\partial r}&\frac{\partial x_{2}}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos(\theta)&-r\sin(\theta)\\ \sin(\theta)&r\cos(\theta)\end{vmatrix}=r.$
-Then,
-$dA=r~dr~d\theta.$
-Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
+Let $\tilde{f}$ be the disintegration map of $\mu$ with respect to $\pi_{*}\mu$.
+We will show that $\mathcal{E}_{1}(\tilde{f})=1$.
+Consider this setup in polar coordinates. Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
 $d\mu=\frac{dA}{A}=\frac{r~dr~d\theta}{\pi R^{2}}.$
-Define $\nu:=\pi_{*}\mu$. Note that, give a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between circles of radii $a$ and $b$, normalized by $A$:
+Define $\nu:=\pi_{*}\mu$. Note that, given a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between circles of radii $a$ and $b$, normalized by $A$:
 $\nu((a,b))=\frac{\pi b^{2}-\pi a^{2}}{\pi R^{2}}=\frac{b^{2}-a^{2}}{R^{2}}.$
@@
 Note that
-$\int_{Y}\mu_{y}~\nu=\int_{0}^{R}\int_{0}^{2\pi}\Big{(}\frac{1}{2\pi}\delta_{y}d\theta\Big{)}\frac{2r}{R^{2}}~dr=\frac{1}{\pi R^{2}}\iint r~dr~d\theta$
+$\int_{Y}\mu_{y}~d\nu=\int_{0}^{R}\int_{0}^{2\pi}\Big(\frac{1}{2\pi}\delta_{y}d\theta\Big)\frac{2r}{R^{2}}~dr=\frac{1}{\pi R^{2}}\iint r~dr~d\theta,$
 so we recover the measure $\mu$ by integrating the family $\mu_{y}$ with respect to $\nu$.
-Moreover, $d_{Y}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=|y-y^{\prime}|$.
-For simplicity, consider $W_{1}$.
+Moreover, $d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=|y-y^{\prime}|$.
 One way to write the 1-Wasserstein distance is
@@
 $W_{1}(\mu_{y},\mu_{y^{\prime}})=\int_{F_{y}}|y-y^{\prime}|~d\mu_{y}=|y-y^{\prime}|.$
-Then, the disintegration map $\tilde{f}$ of $\mu$ with respect to $\nu$ is a isometry, and $\mathcal{E}_{1}(\tilde{f})=1$.
-Similarly, consider the case $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}\leq R^{2}\}$, with $R>0$ and $\lambda=1+\varepsilon$, $\varepsilon$ positive and sufficiently small, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed, $Y=(0,R]$, and the foliation of $X$ is given into ellipses, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+\lambda^{2}x_{2}^{2}}=y$.
-Converting into polar coordinates: consider $x_{1}=r\cos(\theta)$, $x_{2}=\frac{r}{\lambda}\sin(\theta)$, and
-$J(r,\theta)=\begin{vmatrix}\frac{\partial x_{1}}{\partial r}&\frac{\partial x_{1}}{\partial\theta}\\ \frac{\partial x_{2}}{\partial r}&\frac{\partial x_{2}}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos(\theta)&-r\sin(\theta)\\ \frac{1}{\lambda}\sin(\theta)&\frac{r}{\lambda}\cos(\theta)\end{vmatrix}=\frac{r}{\lambda}.$
-Then,
-$dA=\frac{r}{\lambda}~dr~d\theta.$
-Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
+Then, $\mathcal{E}_{1}(\tilde{f})=1$.
+[…]
+Consider $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}\leq R^{2}\}$, $R>0$, $\lambda\geq 1$, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed on $X$, $Y=(0,R]$, and a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+\lambda^{2}x_{2}^{2}}=y$.
+In polar coordinates
 $d\mu=\frac{dA}{A}=\frac{\frac{r}{\lambda}~dr~d\theta}{\frac{\pi R^{2}}{\lambda}}=\frac{r~dr~d\theta}{\pi R^{2}}.$
-Define $\nu:=\pi_{*}\mu$. Note that, give a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between ellipses of major axis $a$ and $b$, normalized by $A$:
+Define $\nu:=\pi_{*}\mu$. Note that, given a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between ellipses of major axis $a$ and $b$, normalized by $A$:
 $\nu((a,b))=\frac{\frac{\pi b^{2}}{\lambda}-\frac{\pi a^{2}}{\lambda}}{\frac{\pi R^{2}}{\lambda}}=\frac{b^{2}-a^{2}}{R^{2}}.$
@@
 However, note that in this case the relation arc length and $\theta$ is no longer linear (see Figure 3).
-Figure 3. Normalized arc length as a function of $\theta$ for $\lambda$ equals to $1,1.1,1.5$ and $2$.
-Therefore, the measures $\mu_{y}$ have a different dependence on $\theta$:
-$d\mu_{y}(r,\theta)=\frac{\frac{dL_{y}(\theta)}{d\theta}}{\mathcal{L}_{y}}~\delta_{y}~d\theta$
-where
-$L_{y}(\theta)=\int_{0}^{\theta}y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\tilde{\theta})}~d\tilde{\theta}$
-is the arc length of the ellipse $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$ for angle between 0 and $\theta$, and $\mathcal{L}_{y}=L_{y}(2\pi)$. By the Fundamental Theorem of Calculus
+Figure 3. Normalized arc length in function of $\theta$ for different $\lambda$.
+We have
 $d\mu_{y}(r,\theta)=\frac{y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\theta)}}{\mathcal{L}_{y}}~\delta_{y}~d\theta.$
-In this case, $d_{Y}(y,y^{\prime})=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\frac{1}{\lambda}|y-y^{\prime}|$ and, to calculate the $1$-Wasserstein distance between $\mu_{y}$ and $\mu_{y^{\prime}}$, we still have the radial transport, although $d(x,T(x))$ depends of $\theta$: for $x=y\cos(\theta)$, $d(x,T(x))=R_{y^{\prime}}(\theta)-R_{y}(\theta)$, where
+In this case, $d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\frac{1}{\lambda}|y-y^{\prime}|$ and, to calculate the $1$-Wasserstein distance between $\mu_{y}$ and $\mu_{y^{\prime}}$, we still have the radial transport, although $d(x,T(x))$ depends of $\theta$: for $x=y\cos(\theta)$, $d(x,T(x))=R_{y^{\prime}}(\theta)-R_{y}(\theta)$, where
 $R_{y}(\theta)=\frac{y}{\sqrt{\lambda^{2}-(\lambda^{2}-1)\cos^{2}(\theta)}}$
@@
 $=$ $\frac{2\pi~(y^{\prime}-y)~y}{\lambda~\mathcal{L}_{y}}.$
-Therefore, considering $\hat{f}_{\lambda}$ the disintegration map of $\mu$ with respect to $\nu$, we have
-$\mathcal{E}_{1}(\hat{f}_{\lambda})=\frac{2\pi y}{\mathcal{L}_{y}},$
+Let $\tilde{f}_{\lambda(t)}$ be the disintegration map of $\mu_{t}$ with respect to $\nu$. Based on the developments above, we have that
+$\mathcal{E}_{1}(\tilde{f}_{\lambda(t)})=\frac{2\pi y}{\mathcal{L}_{y}},$
 that is, the energy is the ratio between the length of the circle with radius $y$ and the length of the ellipse with major axis equal to $y$.
-Table 1 displays the values of $\mathcal{L}_{y}$ for $\lambda=1$ (circle), $\lambda=1,001$, $\lambda=1,01$, $\lambda=1,1$, $\lambda=1,5$ and $\lambda=2$.
+Thus, we have that as $\mu$ evolves over time, the structure for the disintegration of $\mu_{t}$ with respect to $\pi_{*}\mu_{t}$ also changes, and this variation is captured by the energy functional.
+Indeed, Figure 4 depicts the dependence of $\mathcal{E}_{1}(\tilde{f}_{\lambda(t)})$ on $\lambda(t)$.
+As shown, $\mathcal{E}_{1}$ grows smoothly with increasing foliation perturbation.
+Figure 4. Energy as a function of the eccentricity of the leaves.
+For comparison purposes, Table 1 displays the values of $\mathcal{L}_{y}$ for $\lambda=1$ (circle), $\lambda=1,001$, $\lambda=1,01$, $\lambda=1,1$, $\lambda=1,5$ and $\lambda=2$.
 Table 2 summarizes the computed values of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ corresponding to the cases in Table 1.
-Observe that $\mathcal{E}_{1}(\hat{f}_{\lambda})$ is independent of $y$.
 Table 1. $\mathcal{L}_{y}$ for different values of $\lambda$
@@
 11 $1,000499687$ $1,004968906$ $1,047025412$ $1,188089106$ $1,297046785$
-Figure 4 depicts the dependence of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ on $\lambda$. As shown, $\mathcal{E}_{1}$ grows smoothly with increasing foliation perturbation.
-Figure 4. $\mathcal{E}_{1}(\hat{f}_{\lambda})$ as a function of $\lambda$
-This example highlights the versatility of our approach: beyond classifying foliations on $X$ via disintegration maps, it also provides a framework for analyzing structural perturbations in foliations.
-Notably, the concepts developed here have potential applications across dynamical systems, machine learning, and stochastic analysis.
+Example 4.6 highlights the versatility of our approach: beyond characterize foliations on $X$ via disintegration maps, it also provides a framework for analyzing structural perturbations in foliations.
+The concepts developed here have potential applications across Dynamical Systems, Machine Learning, and Stochastic Analysis.
 Acknowledgements
@@
 would like to acknowledge support from the Max Planck Institute for Mathematics in the Sciences and from CAPES.
+Our manuscript has no associated data.
+The authors declare no conflict of interest.